solve the system of equations
y= 2x-4
y= 4x-10

Answer by algebrahouse.com(1659)

y = 2x - 4
y = 4x - 10

Since both expressions are equal to y, then they are equal to each other.

2x - 4 = 4x - 10 {substituted 2x - 4, in for y, into bottom equation}
-4 = 2x - 10 {subtracted 2x from each side}
2x = 6 {added 10 to each side}
x = 3 {divided each side by 2}

y = 2x - 4 {top equation}
y = 2(3) - 4 {substituted 3, in for x, into top equation}
y = 6 - 4 {multiplied}
y = 2 {subtracted}

x = 3 and y = 2

By the way, (3,2) is the point of intersection of the two lines when
they are graphed on a coordinate plane.

For more help from me, visit: www.algebrahouse.com